OpenStudy (anonymous):

Wendell is looking over some data regarding the strength of some building materials and how that relates to their length. The data seems to be exponential. How can Wendell determine if the data is exponential, and how can he convert that to a logarithmic function?

OpenStudy (perl):

Suppose $$y=Ae^{kx}$$ Then, taking logs, we see that $$ \log y=kx+\log A$$. Therefore, try taking logs of the y-values, and plot those against the ordinary x-values. You'll get a straight line whose slope tells you k and whose y-intercept tells you log A. here $$ \log x $$ refers to the natural logarithm ln(x)

OpenStudy (anonymous):

So is that like the whole answer? How do we know if it's exponential.